Lux Autocorp is a company specialized in the production of luxury vehicles. Let’s assume its total cost of producing 4 cars is $450,000 and its total cost of producing 5 cars is $500,000. a) What is the average total cost of producing 4 cars? b) What is the marginal cost of the fifth car?
Answer #1

A.

Average cost of producing 4 cars

=total cost / quantity

=4,50,000/4

=1,12,500.

B.

Marginal cost= change in total cost / change in quantity

=(5,00,000-4,50,000)/(5-4)

=50,000/1

=50,000.